OpenStudy (pokemon23):

3-y=8y 3-y=8y @Mathematics

OpenStudy (anonymous):

y=3

jimthompson5910 (jim_thompson5910):

3-y=8y 3=8y+y 3=9y 3/9 = y 1/3 = y y = 1/3